# Homophily Within and Across Groups

### v2

@@ Abstract
Homophily—the tendency to connect with similar others—plays a central role in shaping network structure and function. It is often treated as a uniform, global parameter, independent from other structural features. Here, we propose a maximum-entropy framework that decomposes homophily into contributions within and across groups, with the stochastic block model emerging as a special case. Our exponential-family formulation, parameterized by group size, fits real-world social networks well and allows homophily to be captured with a single parameter per group size. This decomposition shows that aggregate metrics often obscure group-level variation. We also find that the scale-dependent distribution of homophily has a significant impact on network percolation, influencing epidemic thresholds, the spread of ideas or behaviors, and the effectiveness of intervention strategies. Ignoring this heterogeneity can lead to distorted conclusions about connectivity and dynamics in complex systems.
@@ I Introduction
Homophily—the tendency to interact with similar others—shapes both the structure and function of networks [1, 2, 3, 4, 5]. In social systems, people tend to connect with others who share attributes such as sex, age, race, occupation, sexual preferences, socioeconomic background, or vaccination status [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. Assortative mixing occurs in various settings, ranging from friendships and workplace ties to online platforms, peer influence networks [17, 18, 19, 20], and even graph neural networks [21, 22, 23, 24]. It shapes interactions and governs processes like contagion, diffusion, and coordination [25, 26, 27, 28, 29]. In healthcare settings, the close contact between healthcare workers and vulnerable patients can alter outbreak patterns, as homophily interacts with variations in infectiousness and susceptibility, reshaping epidemic dynamics [30, 31, 32, 33, 34, 35, 36, 37].
@@
Homophily varies by interaction type and tie strength [38, 39, 40]. Strong ties tend to form within close-knit groups, while weak ties more often bridge distant or dissimilar clusters [41, 39]. Networks of citations and trade partnerships typically exhibit dual mixing patterns, combining dense local collaboration with broad cross-group exchange [42, 43]. Gender homophily, in particular, is sensitive to both medium and group size. On Instagram, women show strong same-gender preferences in comments, while men are largely neutral [44]. In face-to-face proximity networks, women tend to form homophilic triads, whereas men favor same-sex dyads [45]. In contrast, the teen-focused Spanish social network Tuenti shows the opposite: women sustain stronger same-gender dyads, and men more often form all-male triads than expected under null models [46].
@@
These observations indicate that homophily is influenced by both group composition and size [47], highlighting its variability across social scales [48]. Small, cohesive groups often display high local homophily, whereas long-range connections mix more randomly, producing distinct within- and across-group connectivity essential for understanding collective behavior [49]. Nevertheless, many network models simplify homophily into a single, uniform parameter—essentially a summary statistic that captures only the network’s average mixing pattern [50, 25, 14, 51, 52, 53, 54]. Epidemiological studies emphasize the importance of this complexity, demonstrating that high within-group homophily can elevate epidemic thresholds, while increased cross-group interactions typically lower them [6, 55, 56, 57, 58, 59]. Thus, to improve our understanding of network formation and dynamics, we must model homophily as a heterogeneous quantity across social scales.
@@
We introduce a network model in which groups, modeled as cliques, exhibit homophily that varies with group size. Using a maximum-entropy framework, the model captures homophilic structure beyond pairwise links and distinguishes between within- and across-group connections. This reveals scale-dependent patterns often overlooked by standard measures and aligns well with empirical studies. Our results show that such multi-scale homophily can raise or lower the percolation threshold, depending on how intra- and inter-group links are distributed. These findings offer new tools for understanding epidemic thresholds, designing interventions, and modeling the spread of information or behavior in social systems.
@@
The remainder of the paper is structured as follows. Section II describes our network model and demonstrates how it encodes multi-scale homophily. Section III validates the model using real-world data, illustrating how within- and across-group mixing patterns emerge under the same modeling assumptions. Section IV examines how multi-scale homophily influences percolation thresholds and network connectivity, with implications for epidemic control and other contagion processes.
@@
Figure 1: Constructing a Homophilic Clique Network with Eight Red and Seven Blue Nodes. With two colors, each $c$-clique can appear in $c+1$ different compositions, distributed according to ${F}_{c}$. (a-c) The network $G$ is constructed by merging multiple clique layers—specifically, 2-, 3-, and 4-cliques—where each layer $G_{c}$ contains cliques of size $c$ but shares the same set of nodes. While the nodes remain fixed, each layer differs in clique size $c$ and composition distribution ${F}_{c}$. The accompanying histograms show the frequency of each $c$-clique type in the network $G_{c}$. In each layer, $M_{c}$ groups of $c$ nodes are sampled with replacement from the available colored nodes and converted into $c$-cliques. For example, $G_{3}$ consists of seven sampled 3-cliques, while $G_{2}$ is constructed similarly, but with its own parameters $M_{2}$ and ${F}_{2}$. (d) All layers are then merged to form the final network: $G=\bigcup G_{c}$. (e) The maximum-entropy clique composition distribution ${F}_{5}$, derived from Eq. (8), is illustrated for cliques of size $c=5$ at homophily levels $h_{5}=0.2$ and $h_{5}=0.6$, across several values of the red-node fraction $n_{\mathrm{r}}$. For clarity, we omit the corresponding networks generated from these distributions. More details in Sec. II.
@@ II Model Description
We introduce a network model that captures homophily at different social scales. The model generates an undirected, sparse graph $G$, formed as the union of multiple subnetworks $\{G_{c}\}$, all sharing the same set of nodes. Each subnetwork $G_{c}$ consists of cliques of size $c$ (groups). Every node in the network has a discrete attribute—for example, a categorical label indicating membership. We begin with a binary labeling of nodes as red or blue, extending to multiple classes in Sec. V.4. As shown in Fig.1, the network consists of $N$ nodes, with $N_{\mathrm{r}}$ red and $N_{\mathrm{b}}=N-N_{\mathrm{r}}$ blue. Their corresponding proportions are $n_{\mathrm{r}}=N_{\mathrm{r}}/N$ and $n_{\mathrm{b}}=1-n_{\mathrm{r}}$. All subnetworks $G_{c}$ share the same node set; that is, red and blue labels are consistent across layers.
We classify the cliques within each subnetwork $G_{c}$ based on the number of red or blue nodes they contain. A $c$-clique can be classified into one of $c+1$ types, that are specified by the number $i$ of red nodes included. We denote by $F_{c,i}$ the proportion of $c$-cliques in $G_{c}$ that contain exactly $i$ red nodes and $c-i$ blue nodes. To ensure that the overall fraction of red and blue nodes is preserved in the large-network limit, the distributions $\{F_{c,i}\}$ must satisfy
@@
To construct each subnetwork $G_{c}$, we independently sample $M_{c}$ cliques of size $c$ from the total population of nodes according to the distribution $\{F_{c,i}\}$. Fig. 1(a-c) illustrates the construction of three subnetworks with different clique-type distributions. In this framework, $G_{1}$ is a trivial graph of isolated colored nodes, and $G_{2}$ is a two-block stochastic block model [60]. The final network $G$, as shown in Fig. 1d, is obtained by merging all subnetworks;
@@
Since cliques are sampled independently, setting $M_{c}=\mathcal{O}(N)$ ensures that each layer $G_{c}$ remains sparse and that the fraction of overlapping links vanishes in the large-network limit. Given a prescribed average degree $\langle d\rangle$, controlling $M_{c}$ allows us to regulate the number of links from each clique size to the overall network. In practice, we sample fewer large cliques, reflecting empirical observations: larger cliques tend to be rarer in real-world networks due to cognitive and social constraints—a phenomenon known as “schisming” [61, 62]. A more detailed discussion regarding finiteness and empirical networks is provided in Sec. V.2.
@@ II.1 Group Homophily Values
where $e_{\mathrm{rr}}$ and $e_{\mathrm{bb}}$ are the fractions of red–red and blue–blue links, and $n_{k}$ is the fraction of nodes of color $k$. The numerator in Eq. 3 corresponds to modularity, while $h$ is modularity normalized by its maximum value [66, 67]. By construction, $|h|\leq 1$: values near 1 signal strong homophily or segregation; $h\approx 0$ indicates random mixing; and negative values imply heterophily [68].
@@ II.2 Maximum Entropy Distribution
This two-color framework easily generalizes to multiple categorical attributes and even multidimensional homophily, allowing more attributes to be considered simultaneously. Sections V.4 and V.5 detail these generalizations explicitly.
@@ III Results
For each empirical network $G$, we extract all maximal cliques and group them by size to form distinct layers $\{G_{c}\}$. As shown in Fig. 1(a–d), this decomposition yields a family of subgraphs, each capturing interactions at a specific clique size. Some links appear in multiple layers, but this redundancy has a negligible effect on measured homophily. We fit the maximum-entropy model (8) to the observed clique-type distributions by maximizing the likelihood, thereby estimating the parameters $\{\theta\}$. For details on edge duplication and its impact, see Sec. V.2.
@@
Total Coleman indices range from mildly homophilic ($h=0.1$ in the Last.fm network and $h=0.2$ in the University of Oklahoma Facebook network), to randomly mixed ($h=0.0$ in the CPH and SocioPatterns networks), and mildly heterophilic ($h=-0.2$ on the Pokec platform). However, these aggregate measures conceal substantial variation at the clique level. We observe diverging trends across datasets. In the mildly homophilic networks, homophily increases with clique size: Last.fm exhibits random mixing for dyads ($h=0$) but stronger homophily in larger cliques ($h=0.3$); Oklahoma shows heterophily among bridging links ($h=-0.2$) and homophily only in larger groups. In contrast, CPH and SocioPatterns—despite appearing randomly mixed overall—display the opposite pattern: bridging links are homophilic, but larger cliques are heterophilic.
The Pokec network offers a particularly striking example of a pattern behind an average index. Its Coleman index $h=-0.2$ suggests mild heterophily, yet dyads ($G_{2}$) exhibit strong cross-sex preference with $h_{2}=-0.6$. As clique size increases from 2 to 10, this intermingling turns into homophily. This transformation is not a minor refinement—it fundamentally changes the interpretation of the network: pairwise links reflect Pokec’s dating-oriented function, while larger cliques reflect its broader role as a social media platform where, at social scales, birds of a feather flock together [1].
Dataset Link Type $n_{\mathrm{r}}$ (%) $h$ $N$
@@
Figure 2: Empirical and Theoretical Distributions of Clique Types and Homophily in Real-world Networks mentioned in Sec. III. (a) Each column compares empirical clique-type distributions (histograms) with the corresponding maximum-entropy distributions (dashed lines, computed using Eq. (8)), showing strong agreement and validating the model’s accuracy in capturing observed clique compositions across real-world datasets. Red represents males and blue, females. (b) Variation in homophily values $h_{c}$ across clique sizes (up to $c=10$) for several empirical networks, highlighting distinct group interaction patterns based on sex attributes. Error bars represent 95% confidence intervals obtained via bootstrap resampling. (c–e) Average homophily trends across clique sizes for Facebook friendship networks from 100 U.S. institutions, with attributes grouped by: (c) Sex (two categories), (d) Student status (two categories), and (e) Class year (12 categories). Gray background curves represent individual institutions, while colored lines indicate the average trends. Lighter colors reflect a greater number of contributing institutions, emphasizing both inter-institutional variability and how homophily depends on the attribute under consideration.
Facebook data from 100 U.S. universities and colleges offers a socially comparable set of environments with rich attribute annotations. Altenburger et al. [82] previously showed that users tend to form highly sex-homogeneous friendship circles—often all-male or all-female—even though these preferences largely cancel out in aggregate, producing weak overall homophily. Fig. 2(c–e) illustrates how homophily varies with clique size when nodes are grouped by (c) sex, (d) student status, and (e) class year. Each attribute captures different social dynamics, leading to distinct homophily profiles. In Fig. 2c, sex-based homophily increases monotonically with clique size, suggesting that sex-based clustering intensifies in larger groups. In contrast, Fig. 2d shows an average trend with a medium level of homophily at the link level and only a slight dependency on group size.
@@
IV Percolation
The structure of the network determines how robust it is to being disintegrated into small, connected components when links or nodes are removed [85, 86, 87, 88, 89]. In social systems, these removals can stem from communication failures, a limited number of contacts, or other reasons, and the connectivity then determines the spread of ideas, behaviors, or diseases [90, 38].
Removing across-group links—often corresponding to weak ties that bridge otherwise separate communities [91]—can trigger sudden fragmentation, severing pathways across the network. In contrast, losing within-group links, typically associated with strong local ties within groups, tends to undermine cohesion within groups without immediately disrupting global connectivity [92, 93, 94]. This distinction is central to our analysis, where we explicitly differentiate between links that connect groups (the $G_{2}$-layer of our model) and those that bind them internally (the $G_{c}$ layers for $c>2$).

### v3

@@ Abstract
Homophily—the tendency of individuals to interact with similar others—shapes how networks form and function. Yet existing approaches typically collapse homophily to a single scale, either one parameter for the whole network or one per community, thereby detaching it from other structural features. Here, we introduce a maximum-entropy random graph model that moves beyond these limits, capturing homophily across all social scales in the network, with parameters for each group size. The framework decomposes homophily into within- and across-group contributions, recovering the stochastic block model as a special case. As an exponential-family model, it fits empirical data and enables inference of group-level variation of homophily that aggregate metrics miss. The group-dependence of homophily substantially impacts network percolation thresholds, altering predictions for epidemic spread, information diffusion, and the effectiveness of interventions. Ignoring such heterogeneity risks systematically misjudging connectivity and dynamics in complex systems.
@@ I Introduction
Homophily—the tendency for like to connect with like—shapes both the structure and function of networks [1, 2, 3, 4, 5]. In social systems, people tend to connect with others who share attributes such as gender, age, race, education, occupation, sexual preferences, socioeconomic background, or vaccination status [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Assortative mixing occurs in various settings, ranging from friendships and workplace ties to online platforms, peer influence networks [18, 19, 20, 21], and even graph neural networks [22, 23, 24, 25]. It shapes interactions and governs processes like contagion, diffusion, and coordination [26, 27, 28, 29, 30]. In healthcare settings, the close contact between healthcare workers and vulnerable patients can alter outbreak patterns, as homophily interacts with variations in infectiousness and susceptibility, reshaping epidemic dynamics [31, 32, 33, 34, 35, 36, 37, 38].
@@
Homophily varies by interaction type and tie strength [39, 40, 41]. Strong ties tend to form within close-knit groups, while weak ties more often bridge distant or dissimilar clusters [42, 40]. Networks of citations and trade partnerships typically exhibit dual mixing patterns, combining dense local collaboration with broad cross-group exchange [43, 44]. Gender homophily, in particular, is sensitive tothe interaction context and size. On Instagram, women show strong same-gender preferences in comments, while men are largely neutral [45]. In face-to-face proximity networks, women tend to form homophilic triads, whereas men favor same-sex dyads [46]. In contrast, the teen-focused Spanish social network Tuenti shows the opposite: women sustain stronger same-gender dyads, and men more often form all-male triads than expected under null models [47].
@@
These observations suggest that homophily is influenced by both group composition and size, underscoring its variability across different social scales [48, 49]. Small, cohesive groups often display high local homophily, whereas long-range connections mix more randomly, producing distinct within- and across-group connectivity essential for understanding collective behavior [50]. Nevertheless, many network models simplify homophily into a single, uniform parameter—essentially a summary statistic that captures only the network’s average mixing pattern [51, 26, 14, 52, 53, 54, 55]. On the other hand, models with explicit community structure, such as the Stochastic Block Model, allow different homophily values across communities but typically assign each node to a single group (and thus a single effective length scale), so they cannot capture overlapping, multiscale affiliations in which a node simultaneously participates in groups of different sizes. Epidemiological studies emphasize the importance of this complexity, demonstrating that high within-group homophily can elevate epidemic thresholds, while increased cross-group interactions typically lower them [6, 56, 57, 58, 59, 60]. Recent studies further show that variations in homophily across interaction sizes critically influence how quickly different groups gain access to information [61]. Therefore, to better understand network formation and dynamics, we should model homophily as a heterogeneous property that varies across social scales.
@@
We introduce a network model that allows for homophily to operate across the network scale and scales of various finite social groups. These groups are operationalized by cliques, which serve as natural building blocks of human interactions. The cliques can represent various social foci [62] and social circles [63, 39], which can consist of families, workplaces, organizations, friendship groups, and other social, physical, legal, or psychological entities. In our framework, we model the homophily within cliques such that the level of homophily depends on the size of the cliques. This allows, for example, our model to have different homophily for interactions within groups and bridging groups (cliques with two nodes). By adopting a maximum-entropy formulation, the model captures homophilic structure beyond pairwise links and distinguishes between within- and across-group connections. This reveals scale-dependent patterns often overlooked by standard measures and aligns well with empirical studies. Our results show that such multi-scale homophily can raise or lower the percolation threshold, depending on how intra- and inter-group links are distributed. These findings offer new tools for understanding epidemic thresholds, designing interventions, and modeling the spread of information or behavior in social systems.
@@
The remainder of the paper is organized as follows. In Sec. II, we present the generative network model and demonstrate how it encodes multi-scale homophily. In Sec. III, we validate the model against empirical data, illustrating how distinct within- and across-group mixing patterns emerge naturally within this unified framework. We then investigate how these multi-scale homophily patterns affect percolation thresholds and network connectivity, highlighting implications for epidemic control strategies and other contagion processes.
@@
Figure 1: Constructing a Homophilic Clique Network with Eight Red and Seven Blue Nodes. With two colors, each $c$-clique can appear in $c+1$ different compositions, distributed according to ${F}_{c}$. (a-c) The network $G$ is constructed by merging multiple clique layers—specifically, 2-, 3-, and 4-cliques—where each layer $G_{c}$ contains cliques of size $c$ but shares the same set of nodes. While the nodes remain fixed, each layer differs in clique size $c$ and composition distribution ${F}_{c}$. The accompanying histograms show the frequency of each $c$-clique type in the network $G_{c}$. In each layer, $M_{c}$ groups of $c$ nodes are sampled with replacement from the available colored nodes and converted into $c$-cliques. For example, $G_{3}$ consists of seven sampled 3-cliques, while $G_{2}$ is constructed similarly, but with its own parameters $M_{2}$ and ${F}_{2}$. (d) All layers are then merged to form the final network: $G=\bigcup G_{c}$. (e) The maximum-entropy clique composition distribution ${F}_{5}$, derived from Eq. (8), is illustrated for cliques of size $c=5$ at homophily levels $h_{5}=0.0,h_{5}=0.2$ and $h_{5}=0.6$, across several values of the red-node fraction $n_{\mathrm{r}}$. For clarity, we omit the corresponding networks generated from these distributions. More details in Sec. II.
@@ II Model Description
We introduce a generative maximum-entropy model that both fits empirical data and provides mechanistic insight, while serving as a statistical framework for inferring group-level variations in homophily. The model is a random graph construction designed to capture homophily across multiple social scales (clique sizes). Rather than enforcing node-specific degree constraints, we fix only the average degree, allowing node degrees to fluctuate around the mean, yielding an approximately Poisson distribution.
We begin with a binary attribute assignment: each of the $N$ nodes is labeled red or blue, with $N_{\mathrm{r}}$ red nodes and $N_{\mathrm{b}}=N-N_{\mathrm{r}}$ blue nodes. Their proportions are given by $n_{\mathrm{r}}=N_{\mathrm{r}}/N$ and $n_{\mathrm{b}}=1-n_{\mathrm{r}}$. All subnetworks $\{G_{c}\}$ are defined on this same node set, so red and blue labels are consistent across layers. (The framework extends to multiple attribute classes in Sec. .5.) We classify the cliques within each subnetwork $G_{c}$ based on the number of red or blue nodes they contain. A $c$-clique can be classified into one of $c+1$ types, that are specified by the number $i$ of red nodes included. We denote by $F_{c,i}$ the proportion of $c$-cliques in $G_{c}$ that contain exactly $i$ red nodes and cic-i blue nodes. To ensure that the overall fraction of red and blue nodes is preserved in the large-network limit, the distributions $\{F_{c,i}\}$ must satisfy
@@
To construct each subnetwork $G_{c}$, we independently sample $M_{c}$ cliques of size $c$ from the total population of nodes according to the distribution $\{F_{c,i}\}$. Fig. 1(a-c) illustrates the construction of three subnetworks with different clique-type distributions. In this framework, $G_{1}$ is a trivial graph of isolated colored nodes, and $G_{2}$ is a two-block stochastic block model [64]. The final network $G$, as shown in Fig. 1d, is obtained by merging all subnetworks $\{G_{c}\}$, where each subnetwork consists of cliques of size $c$ drawn over the same set of nodes;
@@
Since cliques are sampled independently and we require the largest clique size to be finite, setting $M_{c}=\mathcal{O}(N)$ ensures that each layer $G_{c}$ remains sparse and that the fraction of overlapping links vanishes in the large-network limit. For more details, see Methods .2. Given a prescribed average degree $\langle d\rangle$, controlling $M_{c}$ allows us to regulate the number of links from each clique size to the overall network. In practice, we sample fewer large cliques, reflecting empirical observations: larger cliques tend to be rarer in real-world networks due to cognitive and social constraints—a phenomenon known as “schisming” [65, 66]. A more detailed discussion regarding finiteness and the effect of overlap on empirical networks is provided in Methods .3.
@@ II.1 Group Homophily Values
where $e_{\mathrm{rr}}$ and $e_{\mathrm{bb}}$ are the fractions of red–red and blue–blue links, and $n_{k}$ is the fraction of nodes of color $k$. By construction, $|h|\leq 1$: values near 1 signal strong homophily or segregation; $h\approx 0$ indicates random mixing; and negative values imply heterophily [70].
@@ II.2 Maximum Entropy Distribution
The goal of this model is not to replicate every microscopic detail (such as exact degree distributions), but rather to serve as a minimal generative null model that clearly isolates the role of multiscale homophily. This two-color framework naturally generalizes to scenarios with multiple categories or even multidimensional homophily, allowing for the simultaneous consideration of various attributes. We explicitly present these extensions in Sections .5 and .6.
@@ III Results
Here, we first demonstrate that our model accurately captures empirical homophily patterns observed in real-world networks. We then investigate how these biases in connectivity patterns influence dynamical processes unfolding on the networks.
III.1 Model Validation
@@
For each empirical network $G$, we extract all maximal cliques and group them by size to form distinct layers $\{G_{c}\}$. As shown in Fig. 1(a–d), this decomposition yields a family of subgraphs, each capturing interactions at a specific clique size. Some links appear in multiple layers, but this redundancy has a negligible effect on measured homophily. We fit the maximum-entropy model (8) to the observed clique-type distributions by maximizing the likelihood, thereby estimating the parameters $\{\theta\}$. For details on link duplication and its impact, see Sec. .3.
@@
The Coleman indices for these datasets range from mildly homophilic ($h=0.1$ in the Last.fm network and $h=0.2$ in the University of Oklahoma Facebook network), to randomly mixed ($h=0.0$ in the CPH and SocioPatterns networks), and mildly heterophilic (h=0.2h=-0.2 on the Pokec platform). However, these aggregate measures conceal substantial variation at the clique level. We observe diverging trends across datasets. In the mildly homophilic networks, homophily increases with clique size: Last.fm exhibits random mixing for dyads ($h=0$) but stronger homophily in larger cliques ($h=0.3$); Oklahoma shows heterophily among bridging links (h=0.2h=-0.2) and homophily only in larger groups. In contrast, CPH and SocioPatterns—despite appearing randomly mixed overall—display the opposite pattern: bridging links are homophilic, but larger cliques are heterophilic.
The Pokec network offers a particularly striking example of a pattern behind an average index. Its Coleman index h=0.2h=-0.2 suggests mild heterophily, yet dyads ($G_{2}$) exhibit strong cross-sex preference with $h_{2}=-0.6$. As clique size increases from 2 to 10, this intermingling turns into homophily. This transformation is not a minor refinement—it fundamentally changes the interpretation of the network: pairwise links reflect Pokec’s dating-oriented function, while larger cliques reflect its broader role as a social media platform where, at social scales, birds of a feather flock together [1].
Figure 2: Empirical and Theoretical Distributions of Clique Types and Homophily in Real-world Networks mentioned in Sec. III.1. (a) Each column compares empirical clique-type distributions (histograms) with the corresponding maximum-entropy distributions (dashed lines, computed using Eq. (8)), showing strong agreement and validating the model’s accuracy in capturing observed clique compositions across real-world datasets. Red represents males and blue, females. (b) Variation in homophily values $h_{c}$ across clique sizes (up to $c=10$) for several empirical networks, highlighting distinct group interaction patterns based on sex attributes. Error bars represent 95% confidence intervals obtained via bootstrap resampling. (c–e) Average homophily trends across clique sizes for Facebook friendship networks from 100 U.S. institutions, with attributes grouped by: (c) Sex (two categories), (d) Student status (two categories), and (e) Class year (12 categories). Gray background curves represent individual institutions, while colored lines indicate the average trends. Lighter colors reflect a greater number of contributing institutions, emphasizing both inter-institutional variability and how homophily depends on the attribute under consideration.
Dataset Link Type $n_{\mathrm{r}}$ (%) $h$ Size
@@
Facebook data from 100 U.S. universities and colleges offers a socially comparable set of environments with rich attribute annotations. Altenburger et al. [84] previously showed that users tend to form highly sex-homogeneous friendship circles—often all-male or all-female—even though these preferences largely cancel out in aggregate, producing weak overall homophily. Fig. 2(c–e) illustrates how homophily varies with clique size when nodes are grouped by (c) sex, (d) student status, and (e) class year. Each attribute captures distinct social dynamics, resulting in different homophily profiles. In Fig. 2c, sex-based homophily increases monotonically with clique size, suggesting that sex-based clustering intensifies in larger groups. In contrast, Fig. 2d shows an average trend with a medium level of homophily at the link level and only a slight dependency on group size.
@@
III.2 Color Percolation
The structure of the network determines how robust it is to being disintegrated into small, connected components when links or nodes are removed [87, 88, 89, 90, 91]. In social systems, these removals can stem from communication failures, a limited number of contacts, or other reasons, and the connectivity then determines the spread of ideas, behaviors, or diseases [92, 39]. Removing across-group links—often corresponding to weak ties that bridge otherwise separate communities [93]—can trigger sudden fragmentation, severing pathways across the network. In contrast, losing within-group links, typically associated with strong local ties within groups, tends to undermine cohesion within groups without immediately disrupting global connectivity [94, 95, 96]. This distinction is central to our analysis, where we explicitly differentiate between links that connect groups (the $G_{2}$-layer of our model) and those that bind them internally (the $G_{c}$ layers for $c>2$).